# Has your pretrained model improved? a multi-head posterior based approach

### 02 Jan 2024 v1

@@ 4 Experimental Analysis
In this experimental session, we initially carry out experiments on a synthetic dataset to verify the effectiveness of our proposed algorithm. Following this, we further evaluate the results in three areas: the MovieLens dataset (Harper & Konstan, 2015) for relational models, spatial datasets (Gurnee & Tegmark, 2023) for large language models, and the Robustness library (Engstrom et al., 2019a; b; Santurkar et al., 2019; 2020) for image models.
@@ 4.1 Synthetic Dataset: Gaussian Mixture Model (GMM) of Ten Gaussian Distributions
To validate the efficacy of our propose posterior-based embedding evaluation metric, as outlined in Eq. (2), we designe an experiment that encompass three scenarios, each with $10$ clusters. These clusters are either perfectly separated, partially overlapping, or perfectly overlapping, and are all generated using a Gaussian Mixture Model. Figure 1 presents the results from these scenarios. As anticipated, the Average of Log Posterior (ALP) scores for the 10 perfectly separated Gaussian Distributions was 00, and the accuracy of the clusters assigned from the posterior matrix was $100\%$. In the case of 10 partially overlapping Gaussian Distributions, the ALP score was $-0.3285$, and the accuracy of the clusters assigned from the posterior matrix was $86.96\%$. Similarly, for the 10 perfectly overlapping Gaussian Distributions, the ALP score was $-0.9372$, and the accuracy of cluster assignment was $57.34\%$.
@@
Figure 1: Illustration on $2d$ synthetic dataset consisting of 10 gaussian distributions that are perfectly separated, partially overlapping, and perfectly overlapping
@@ 4.2 Moive Lens Dataset for Relational
MovieLens-25M consists of 25,000,095 reviews made by 162,541 reviewers on 59,047 movies. We compare with following models to generate the movie embeddings. Word2vec (Mikolov et al., 2013), PDT (Dai et al., 2023) and SASRec (Kang & McAuley, 2018). From the Word2Vec model, we generate two distinct types of embedding representations, specifically w2v_single and w2v_combine. In the case of w2v_single, we create a sequence of movies for each reviewer’s review, sort it based on time, and then employ the Word2vec model to learn the movie embeddings. On the other hand, w2v_combine not only includes the sequences of movies for reviewers but also incorporates sequences of reviewers for each movie and reviewer/movie pairs as part of the training data. For both SASRec and PDT, we generate sequences of nine movies for each reviewer. Additionally, for PDT we also generate sequences of reviewers for each movie, as this is a necessary input for PDT. SASRec is trained using BPR loss, while PDT is trained using two contrastive losses. Both PDT and SASRec are contextualized embeddings, while w2v_single and w2v_combine are static embeddings. We employ two clustering techniques. The first approach involved clustering by single Meta features, such as year and genre. We also apply Embedding tree based method to generate the tree for both year and genre features and use the leaf nodes as clusters.
@@ 4.2.1 Movie Lens Dataset: Clustering by Year
In this section, we have evaluated and compared four kinds of embedding representations that were trained on the MovieLens Dataset. These were trained at various iteration levels, we use the “year” feature as labels to cluster the embeddings. As illustrated in Figure 2(a), the PDT and SRSREC embedding performed better than all embeddings across all iterations, as seen in the mean of average of log posteriors plot. During the early stages of training from iteration 1 to 16, w2v_combine outperformed w2v_single. However, in the later stages from iteration 16 to 50, w2v_single superseded w2v_combine.
(a) Mean of average of log posteriors
@@
Figure 2: Mean of average of the log posterior and accuracy on the MovieLens dataset by clustering on year.
@@ 4.2.2 Movie Lens Dataset: Clustering by Genre
By using the genre feature, we create clusters with the genre features as labels. We then compute and report the mean of average of log posteriors and the mean of accuracy of assigned clusters. These findings are presented in Figure 3. Contrary to the consistent pattern observed with year features as labels, the genre features do not exhibit a similar consistency. From Figure 3(a), it’s noticeable that the PDT embedding generally outperforms both SASRec and w2v_single over all iterations. Furthermore, SASRec surpasses w2v_single from the 1st to the 40th iteration, after which they both plateau with similar scores. Between the 12th and 36th iterations, w2v_combine is observed to outperform PDT. Moving to the mean accuracy of assigned clusters plot (Figure 3(b)), it’s evident that PDT consistently outperforms all other embedding types across all iterations. Generally, w2v_combine surpasses both SASRec and w2v_single, except for the first and third iterations where SASRec exceeds w2v_combine.
(a) Mean of average of log posteriors
@@
Figure 3: Sum of the log posterior on the MovieLens dataset by clustering on genre.
@@ 4.2.3 Movie Lens Dataset: Clustering with Tree Leaf Nodes on Genre and Year as the Meta Features
We explore the case where we use the tree leaf nodes from the embedding tree that is constructed with both year and genre as meta features. The mean average of log posteriors of assigned clusters are calculated and reported, as shown in Figure 4(a). The PDT and SASRec embeddings consistently surpass other embeddings throughout all iterations. However, we notice that w2v_combine surpasses w2v_single from the 1st to the 26th iteration, but w2v_single overtakes w2v_combine from the 26th to the 50th iteration. The mean accuracy of assigned clusters, illustrated in Figure 4(b), clearly shows that the PDT and SASRec embedding exhibit a steady and consistent increase in performance compared to all other embeddings across all iterations. This is followed by w2v_single, which generally surpasses w2v_combine.
(a) Mean of average of log Posteriors
@@
Figure 4: Sum of the log posterior on the movie lens dataset by clustering with tree leaf nodes.
@@ 4.3 Experiments with Embedding from Llama-2 Models
Datasets: We use the Llama-2 models to generate embeddings for three spatial datasets introduced by (Gurnee & Tegmark, 2023). The details of these datasets are shown in Table 2. Each instance of the datasets is a word/phrase for a location (e.g., the name of city, a university, or a place of interest) and based on the spatial location as meta feature, we generate the clusters for each dataset, i.e., based on the continent, state, and borough of the instances from the World Place, USA Place, and NYC Place dataset, respectively.
@@
Embedding Quality: We feed the three datasets into the three Llama-2 models and get their activations/embeddings at each layer of the models. Specifically, we obtain $32$, $40$, and $80$ sets of embeddings for the three sets of models (as the three models have $32$, $40$ and $80$ layers, respectively). For each set of embeddings, we use our proposed method to compute the posterior of individual instance falling into the correct cluster.
@@
In Fig. 5 and 6, In this section, we showcase the quality of the embeddings across three datasets and models. Each plot’s x-axis represents the percentage of layers pertaining to the respective Llama-2 models, while the y-axis signifies the mean of average log posteriors and accuracy. The results reveal a noticeable upward trend in embedding quality from the initial to the final layers. In the USA_Place and NYC_Place datasets, the green lines, representing the larger model (Llama-2-70b), achieve the highest posteriors and accuracy. This suggests that the Llama-2-70b model encodes location information more effectively than the other models. This conclusion holds true for the initial and final interactions of the World_Place dataset, but the consistency wavers during the middle interactions.
@@
Figure 5: Embedding quality over model layers (average posterior). The dimensions are divided into subsets, each comprising 128 dimensions.
@@
Figure 6: Embedding quality over model layers (accuracy). The dimensions are divided into subsets, each comprising 128 dimensions.
Figure 7: Embedding quality over model layers (average posterior). The dimensions are divided into subsets, each comprising 128 dimensions. The covariance matrix for each subset is diagonalized for regularization.
Figure 8: Embedding quality over model layers (accuracy). The dimensions are divided into subsets, each comprising 128 dimensions. The covariance matrix for each subset is diagonalized for regularization.
5 Experiment analysis of CLIP Embeddings on the BREEDS hierarchy
@@
In this section we assess the ability of our metric to predict the classification performance on embeddings of 3 pretrained CLIP (Radford et al., 2021) models on 4 datasets from the Breeds Hierarchy (Santurkar et al., 2020), namely entity 13, entity 30, living 17 and non living 26.
5.1 Models
@@
For the purpose of this study we use CLIP-ViT transformers ViT-L/14, ViT-B/32, ViT-B/16 trained on 224 px $\times$ 224 px images as image encoders. Where ViT-L/14 encodes each image into a 768 dimensional embedding and ViT-B/16 and ViT-B/32 encode each image as a 512 dimensional embedding. For each of the networks we obtained model embeddings for the final image encoder as well as the mean embeddings of internal Multi-Head Self Attention Layers. We train a linear probe on the ImageNet (Deng et al., 2009) training subset of the Breeds dataset as well learn the parameters for estimating the MAP. We then correlate the performance of these 2 metrics to better understand the behaviour of our trained posterior and present the results in Figure 9.
5.2 Experimental Analysis
Table 3: In this table we show the Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(a) on the training and validation sets of entity 13 dataset of the Breeds Hierarchy.
@@ 5.2 Experimental Analysis
Table 4: Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(b) on the training and validation sets of entity 30 dataset of the Breeds Hierarchy.
Table 5: Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(c) on the training and validation sets of living 17 dataset of the Breeds Hierarchy.
Entity 13
Data $\rightarrow$ Train Set Val. Set
Reg. $\downarrow$ B16 B32 L14 B16 B32 L14
@@
Entity 30
Data $\rightarrow$ Train Set Val. Set
Reg. $\downarrow$ B16 B32 L14 B16 B32 L14
@@
Living 17
Data $\rightarrow$ Train Set Val. Set
Reg. $\downarrow$ B16 B32 L14 B16 B32 L14
@@
Non Living 26
Data $\rightarrow$ Train Set Val. Set
Reg. $\downarrow$ B16 B32 L14 B16 B32 L14
@@
Table 3: In this table we show the Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(a) on the training and validation sets of entity 13 dataset of the Breeds Hierarchy.
Table 4: Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(b) on the training and validation sets of entity 30 dataset of the Breeds Hierarchy.
Table 5: Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(c) on the training and validation sets of living 17 dataset of the Breeds Hierarchy.
Table 6: Pearson’s Correlation between layerwise MAP computed over various regularizations and layerwise linear probe accuracy for 8(d) on the training and validation sets of non living 26 dataset of the Breeds Hierarchy.
@@
For the purpose of this study we measure the correlation between MAP and linear probe accuracies learned and computed over the embeddings of the training set and validation set. The results for which are shown in Table 6, Table 6,Table 6 and Table 6 for 8(a), 8(b), 8(c) and 8(d) respectively. In Figure 9 we demonstrate the on a macro level both MAP and linear probe performance correlate. This is across various settings of regularizations for our learned metric over various Breeds Datasets and CLIP Models.
(a) CLIP Models analyzed with multiple regularizations - entity13. The pearson correlations between layerwise MAPs for various regularizations and layerwise linear probe accuracies is shown in Table 6.
(b) CLIP Models analyzed with multiple regularizations - entity30. The pearson correlations between layerwise MAPs for various regularizations and layerwise linear probe accuracies is shown in Table 6
(c) CLIP Models analyzed with multiple regularizations - living17. The pearson correlations between layerwise MAPs for various regularizations and layerwise linear probe accuracies is shown in Table 6
(d) CLIP Models analyzed with multiple regularizations - nonliving26. The pearson correlations between layerwise MAPs for various regularizations and layerwise linear probe accuracies is shown in Table 6
@@
Figure 9: In this figure we present the evolution of layerwise MAP and linear probe accuracies for all CLIP Models on all datasets across varying regularizations strengths to demonstrate correlations between MAP and linear probe performance. The quantitative results for each settings are shown from Table 6-Table 6.
However despite this correlation, the metric is not a reliable indicator of embedding performance as many a times it can predict lower values for models which have higher linear separability as indicated by the linear probe accuracy for the embedding of that given layer, we show one such counter study in Figure 10 with additional results available in Appendix A. In 9(a) for a MAP model which assumes uncorrelated dimensions we observe that for many earlier and later layers of the CLIP models, the order of performance predicted by log posterior isn’t the same as one predict by linear probes on the embedding of that layer. This is particularly evident in case of ViT-L/14 model at both the initial and final layers of the network. A similar phenomena is also observed in 9(b), which imposes a higher regularization on the class covariance matrix learned to estimate the posterior. But this issue doesn’t completely resolve, although is less severe, when reducing the regularization on the learned class covariance as shown in 9(c) and 9(d) for regularizations $10^{-6}$ and $10^{-9}$ respectively.
(a) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 13 breeds dataset. Class covariance matrix assumes uncorrelated embedding dimensions, thereby just using the diagonal of the covariance matrix.
(b) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 13 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-3}$
(c) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 13 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-6}$
(d) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 13 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-9}$
Figure 10: Layerwise multi model analysis of log posterior performance with linear probe performance for a given regularization parameter over entity 13 Breeds Dataset.
5.3 BREEDS Datasets
Breeds Hierarchy (Santurkar et al., 2020) was by leveraging and modifying the WordNet (Miller, 1994) hierarchy for the ImageNet (Deng et al., 2009) dataset to group together semantically similar classes into 1 superclass. The original purpose of creating this hierarchy was to use the subpopulations present in superclasses to detect a model’s robustness to distribution shift but for the purpose of this study we leverage the entire datasets to evaluate the performance of our metric in predicting generalization of CLIP models on those embeddings.

### 14 Feb 2024 v4

@@ 5 Experimental Analysis
In this experimental session, we initially conducted experiments on a synthetic dataset to verify the effectiveness of our proposed algorithm. Following this, we further evaluate the results in three areas: the MovieLens dataset (Harper & Konstan, 2015) for relational models, spatial datasets (Gurnee & Tegmark, 2023) for large language models, and the Robustness library (Engstrom et al., 2019a; b; Santurkar et al., 2019; 2020) for image models.
@@ 5.1 Synthetic Dataset: Gaussian Mixture Model (GMM) of Ten Gaussian Distributions
To validate the efficacy of our proposed posterior-based embedding evaluation metric, as outlined in Equation 2, we designed an experiment encompassing three scenarios, each with $10$ clusters. These clusters were generated such that they are either perfectly separated, partially overlapping, or perfectly overlapping and are all generated using a Gaussian Mixture Model. Figure 1 presents the results from these scenarios. As anticipated, the Average of the Log Posterior (ALP) scores for the ten (10) perfectly separated Gaussian Distributions was 00, and the accuracy of the clusters assigned from the posterior matrix was $100\%$. In the case of 10 partially overlapping Gaussian Distributions, the ALP score was $-0.3285$, and the accuracy of the clusters assigned from the posterior matrix was $86.96\%$. Similarly, for the ten (10) perfectly overlapping Gaussian Distributions, the ALP score was $-0.9372$, and the accuracy of cluster assignment was $57.34\%$.
@@
Figure 1: Illustration on a 2D synthetic dataset consisting of 10 Gaussian distributions that are perfectly separated, partially overlapping, and perfectly overlapping.
@@ 5.2 Moive Lens Dataset for Relational
MovieLens-25M consists of 25,000,095 reviews made by 162,541 reviewers on 59,047 movies. We compare the following models to generate the movie embeddings. Word2vec (Mikolov et al., 2013), PDT (Dai et al., 2023) and SASRec (Kang & McAuley, 2018). From the Word2Vec model, we generate two distinct types of embedding representations, specifically w2v_single and w2v_combine. In the case of w2v_single, we create a sequence of movies for each reviewer’s review, sort it based on time, and then employ the Word2vec model to learn the movie embeddings. On the other hand, w2v_combine not only includes the sequences of movies for reviewers but also incorporates sequences of reviewers for each movie and reviewer/movie pairs as part of the training data. For both SASRec and PDT, we generate sequences of nine movies for each reviewer. Additionally, for PDT, we generate sequences of reviewers for each movie, as this is a necessary input for PDT. SASRec is trained using BPR loss, while PDT is trained using two contrastive losses. Both PDT and SASRec are contextualized embeddings, while w2v_single and w2v_combine are static embeddings. We employ two clustering techniques. The first approach involves clustering by single meta-features, such as year and genre. We also apply the Embedding tree-based method to generate the tree for both year and genre features and use the leaf nodes as clusters.
@@ 5.2.1 Movie Lens Dataset: Clustering by Year
We evaluated and compared four kinds of embedding representations trained on the MovieLens Dataset. These were trained at various iteration levels; we use the “year” feature as labels to cluster the embeddings. As illustrated Figure 2 (a), the PDT and SRSREC embedding performed better than all embeddings across all iterations, as seen in the Mean of the average of log posteriors plot. During the early stages of training from iteration 1 to 16, w2v_combine outperformed w2v_single. However, in the later stages from iteration 16 to 50, w2v_single superseded w2v_combine.
(a) Mean of the average of log posteriors
@@
Figure 2: Mean of the average of the log posterior and accuracy on the MovieLens dataset by clustering on year.
@@ 5.2.2 Movie Lens Dataset: Clustering by Genre
Here, we created clusters with the genre features as labels. We then compute and report the Mean of the average of log posteriors and the Mean of accuracy of assigned clusters. These findings are presented in Figure 3. Contrary to the consistent pattern observed with year features as labels, the genre features do not exhibit a similar consistency. From Figure 3 (a), it’s noticeable that the PDT embedding generally outperforms both SASRec and w2v_single over all iterations. Furthermore, SASRec surpasses w2v_single from the 1st to the 40th iteration, after which they both plateau with similar scores. Between the 12th and 36th iterations, w2v_combine is observed to outperform PDT. Moving to the Mean accuracy of the assigned clusters plot (Figure 3 (b)), it’s evident that PDT consistently outperforms all other embedding types across all iterations. Generally, w2v_combine surpasses both SASRec and w2v_single, except for the first and third iterations where SASRec exceeds w2v_combine.
(a) Mean of the average of log posteriors
@@
Figure 3: Mean of the average of the log posterior and accuracy on the MovieLens dataset by clustering on the genre.
@@ 5.2.3 Movie Lens Dataset: Clustering with Tree Leaf Nodes on Genre and Year as the Meta Features
We also explored the case where we used the tree leaf nodes from the embedding tree constructed with year and genre as meta-features. The Mean average of log posteriors of assigned clusters is calculated and reported, as shown in Figure 4 (a). The PDT and SASRec embeddings consistently surpass other embeddings throughout all iterations. However, we notice that w2v_combine surpasses w2v_single from the 1st to the 26th iteration, but w2v_single overtakes w2v_combine from the 26th to the 50th iteration. The Mean accuracy of assigned clusters, illustrated in Figure 4 (b), clearly shows that the PDT and SASRec embedding exhibit a steady and consistent increase in performance compared to all other embeddings across all iterations. This is followed by w2v_single, which generally surpasses w2v_combine.
(a) Mean of the average of log posteriors
@@
Figure 4: Mean of the average of the log posterior and accuracy on the MovieLens dataset by clustering with tree leaf nodes.
@@ 5.3 Experiments with Embedding from Llama-2 Models
Datasets: We use the Llama-2 models to generate embeddings for three spatial datasets introduced by (Gurnee & Tegmark, 2023). The details of these datasets are shown in Table 2. Each dataset instance is a word/phrase for a location (e.g., the name of a city, a university, or a place of interest). Based on the spatial location as a meta feature, we generate the clusters for each dataset, i.e., based on the continent, state, and borough of the instances from the World Place, USA Place, and NYC Place dataset, respectively.
@@
Embedding Quality: We feed the three datasets into the three (3) Llama-2 models and get their activations/embeddings at each layer of the models. Specifically, we obtain $32$, $40$, and $80$ sets of embeddings for the three sets of models (as the three models have $32$, $40$, and $80$ layers, respectively). We use our proposed method for each set of embeddings to compute the posterior of individual instances falling into the correct cluster.
@@
From Figure 5 and Figure 6, we present the effectiveness of embeddings across three different datasets and models. The x-axis of each graph indicates the percentage of layers relevant to the corresponding Llama-2 models, while the y-axis represents the Mean average of log posteriors and accuracy. The results show a noticeable upward trend in the quality of embeddings from the initial to the final layers. Specifically, in the World_Place, USA_Place, and NYC_Place datasets, the green lines, which denote the larger Llama-2-70b model, exhibit the highest levels of posteriors and accuracy. This indicates that the Llama-2-70b model is more adept at incorporating location data compared to the other models.
@@
Figure 5: Embedding quality over model layers (average of the log posterior). The dimensions are divided into subsets, each comprising 128 dimensions.
@@
Figure 6: Embedding quality over model layers (accuracy of assigned clusters). The dimensions are divided into subsets, each comprising 128 dimensions.
5.4 Experiment analysis of CLIP Embeddings on the BREEDS hierarchy
@@
In this section, we assess the ability of our metric to predict the classification performance on embeddings of 3 pre-trained CLIP (Radford et al., 2021) models on datasets from the Breeds Hierarchy (Santurkar et al., 2020), namely datasets entity 13, entity 30, living 17 and nonliving 26. Next, we briefly describe the Breeds Dataset and the CLIP models used for our evaluation.
Dataset: Breeds Hierarchy (Santurkar et al., 2020) was created by leveraging and modifying the WordNet (Miller, 1994) hierarchy for the ImageNet (Deng et al., 2009) dataset to group together semantically similar classes into one (1) superclass. The original purpose of creating this hierarchy was to use the subpopulations present in superclasses to detect a model’s robustness to distribution shift. For this experiment, we leverage the entire dataset to evaluate the performance of our metric in predicting the generalization of CLIP models on those embeddings.
@@
Models: For this study, we use CLIP-ViT transformers ViT-L/14, ViT-B/32, and ViT-B/16 trained on 224 px $\times$ 224 px images as image encoders, where ViT-L/14 encodes each image into a 768-dimensional embedding and ViT-B/16 and ViT-B/32 encode each image as a 512-dimensional embedding. We obtained model embeddings for each network from the final image encoder and the mean embeddings of internal multi-head self-attention layers. We train a linear probe on the ImageNet (Deng et al., 2009) training subset of the Breeds dataset and learn the parameters for estimating the log posterior (ALP). To compute the average log posterior, we don’t split the embeddings into multiple blocks; therefore, the average log posterior (ALP) demonstrates the flexibility of our approach. We correlate the performance of these two metrics to understand our trained posterior’s behavior better and present the analysis and results next.
@@
Table 3: Pearson’s Correlation on Breeds Datasets (Part 1): Comparing layerwise log posterior and linear probe accuracy across regularization levels in the training and validation sets of the Breeds Hierarchy.
Entity 13 Dataset
Entity 30 Dataset
Train Set Val. Set
Reg. B16 B32 L14 B16 B32 L14
@@
Train Set Val. Set
Reg. B16 B32 L14 B16 B32 L14
@@
Entity 30 Dataset
Table 4: Pearson’s Correlation on Breeds Datasets (Part 2): Comparing layerwise log posterior and linear probe accuracy across regularization levels in the training and validation sets of the Breeds Hierarchy.
Living 17 Dataset
Non Living 26 Dataset
Train Set Val. Set
Reg. B16 B32 L14 B16 B32 L14
@@
Train Set Val. Set
Reg. B16 B32 L14 B16 B32 L14
@@
Non Living 26 Dataset
@@
For this experiment, we measured the correlation between average log posteriors and linear probe accuracy learned and computed over the Breeds training and validation set embeddings. The results are shown in Table 3 and Table 4 for respective datasets from Figure 7 and Figure 8. Based on those results, we demonstrate that average log posterior and linear probe performance correlates strongly across the depth of the network when measured via Pearson’s correlation. This is across various settings of regularizations (both with Independence assumptions and Tikhonov [^1] https://web.eecs.umich.edu/ aey/recent/regular.pdf Regularization) of the class-wise covariance matrices for our learned average log posterior metric for various Breeds Datasets and CLIP Models.
(a) CLIP Model on entity13.
(b) CLIP Model on entity30.
@@
Figure 7: Evolution of layerwise log posterior and linear probe accuracies for CLIP Models across varying regularization strengths, demonstrating correlations between log posterior and linear probe performance across the depth of various CLIP Models. Quantitative results are shown in Table 3 and Table 4.
(a) CLIP Models on living17.
(b) CLIP Models on nonliving26.
Figure 8: Evolution of layerwise log posterior and linear probe accuracies for CLIP Models across varying regularization strengths, demonstrating correlations between log posterior and linear probe performance across the depth of various CLIP Models. A more detailed breakdown of the results comparing the correlations on each dataset for various settings of regularizations corresponding to Figure 7 and Figure 8 is shown in Figure 9 - Figure 16 of Appendix A.1.
Our results with a layerwise analysis of Pre-trained CLIP Models comparing our metric with a linear probe on internal activations help us assert that the log posterior is predictive of an embedding’s downstream classification performance, even if a set of data that was used to generate the embedding wasn’t part of the models training. With Pearson correlations greater than 0.9 for a majority of settings in Table 3 and Table 4, we can confidently establish its value as a metric that can distinguish between good and bad-performing embeddings generated by various models. Additionally we also show the corresponding spearman correlations in Table 5 and Table 6.
Table 5: Spearman’s Correlation on Entity Datasets
Entity 13 Dataset
Entity 30 Dataset
Train Set Val. Set
Reg. B16 B32 L14 B16 B32 L14
$Diag$98 .99 .95 .99 .99 .94
$10^{-3}$97 .99 .98 .97 .98 .96
$10^{-6}$99 .97 .99 .99 .99 .99
$10^{-9}$99 .98 .98 .99 .99 .99
Train Set Val. Set
Reg. B16 B32 L14 B16 B32 L14
$Diag$99 .99 .98 1.00 .99 .98
$10^{-3}$98 1.0 .99 .97 .98 .96
$10^{-6}$82 .79 .96 .99 1.00 .99
$10^{-9}$72 .67 .91 .99 1.00 .99
Entity 30 Dataset
Table 6: Spearman’s Correlation on Living and Non Living Datasets
Living 17 Dataset
Non Living 26 Dataset
Train Set Val. Set
Reg. B16 B32 L14 B16 B32 L14
$Diag$ 1.00 .99 .99 1.00 .99 .98
$10^{-3}$9 .97 .98 1.00 .99 .99
$10^{-6}$47 .41 .42 .99 1.00 .99
$10^{-9}$3 .16 .26 1.00 1.00 .99
Train Set Val. Set
Reg. B16 B32 L14 B16 B32 L14
$Diag$99 1.00 .97 1.00 1.00 .97
$10^{-3}$99 .99 .98 .97 .97 .95
$10^{-6}$62 .48 .77 .99 .98 .99
$10^{-9}$46 .26 .61 .99 .98 .99
Non Living 26 Dataset